# Making Images Real Again: A Comprehensive Survey on Deep Image Composition

### v7

@@ IV-A Rendering based Methods
Conventional image relighting [138, 162, 203, 242, 181] aims to adjust the appearance of an image or the object in an image as lit by novel illumination. With some adaptation, image relighting can also be used to adjust the foreground appearance according to the illumination of a new background [138, 242, 12, 65], which bears some resemblance to image harmonization. However, they usually achieve this goal by inferring explicit illumination condition, material properties, or 3D geometry, in which the supervision for these information is difficult and expensive to acquire. Besides, they generally have strong assumption for the light source, which may not generalize well to complicated real-world scenes.
@@ IV-B Non-rendering based Methods
Early traditional image harmonization methods [221, 170, 85, 165] performed color transformation for the foreground to match the low-level color statistics between foreground and background. The difference between different methods mainly lies in the matching details. For example, Xue et al. [221] proposed to train a classifier to predict the zone (e.g., low, middle, high) of the histogram which can be best matched between foreground and background, and then adjust the foreground color to match the selected zone between foreground and background. [170] explored decomposing an image into a multi-resolution pyramid with multiple subbands, and performing histogram matching for each subband between foreground and background. Lalonde and Efros [85] proposed to represent foreground and background with color clusters, followed by matching foreground and background color clusters. Song et al. [165] proposed to calculate the color transformation (channel-wise scales) based on gray pixels of foreground and background, because normalized illumination color can be directly derived from the pixel values of gray pixels. In some works on sky replacement [263, 176, 184], they attempted to match the color statistics (e.g., mean, variance) between sky region and non-sky region. Broadly speaking, traditional color transfer methods [148, 214, 41, 144, 39, 2] can also be used for color matching between foreground and background to produce a harmonized image.
@@
With the emergence of image harmonization datasets consisting of paired training data (see Section IV-E), abundant image harmonization methods [185, 25, 27, 116, 122, 173, 18, 226, 202, 142, 190] using paired supervision have been developed. Tsai et al. [185] proposed the first end-to-end CNN network for image harmonization and leveraged auxiliary semantic segmentation branch to enhance the basic image harmonization network. Another work Sofiiuk et al. [164] also utilized high-level semantic features, which are inserted into the encoder to provide auxiliary information. Cun and Pun [29] designed an additional Spatial-Separated Attention Module to deal with foreground and background feature maps separately. Hao et al. [58] employed self-attention [198] mechanism to propagate relevant information from background to foreground. By treating different capture conditions as different domains, Cong et al. [25] proposed a domain verification discriminator to pull close the foreground domain and background domain. Similarly, Cong et al. [26] formulated image harmonization as background-guided domain translation task, in which the domain code of background is directly used to guide the harmonization process. One byproduct of [26] is predicting the inharmony level of an image by comparing the domain codes of foreground and background, so that we can selectively harmonize those apparently inharmonious composite images. Inspired by [26], Valanarasu et al. [186] proposed to extract style code from part of background.
@@
Some works [134, 157, 16] concurred that dynamic kernels acting upon feature maps can boost the harmonization performance. Furthermore, [134, 157] pointed out the importance of global information in dynamic kernel prediction. Niu et al. [133] studied domain adaptive image harmonization by treating different datasets as different domains. Specifically, an automatic augmentation network was developed to enrich the illumination diversity of a target domain with limited data.
@@ IV-C Variants of Image Harmonization Task
Subsequently, some works [103, 104, 212, 211, 247, 161] focused on inharmonious region localization task, which aims to localize the inharmonious region in an image. Liang et al. [103] explored aggregating multi-scale contextual information and suppressing redundant information. The methods [104, 212] proposed to magnify the domain discrepancy between foreground and background using color mapping for ease of identifying the inharmonious region.
@@ IV-D Related Research Fields
Image harmonization is closely related to style transfer. Note that both artistic style transfer [46, 68, 140] and photorealistic style transfer [123, 101] belong to style transfer. Image harmonization is closer to photorealistic style transfer, which transfers the style of a reference photo to another input photo. There are two main differences between image harmonization and photorealistic style transfer. 1) Firstly, image harmonization adjusts the foreground appearance according to the background, which needs to take the foreground location into consideration due to the locality property. In contrast, photorealistic style transfer adjusts the appearance of a whole input image according to another whole reference image. 2) Secondly, the definition of “style” in photorealistic style transfer is unclear and coarsely depends on the employed style loss (e.g., Gram matrix loss [46], AdaIn loss [68]). Differently, the goal of image harmonization is clearly adjusting the illumination statistics of foreground, so that the resultant foreground looks like the same object captured in the background illumination condition.
@@ IV-E Datasets and Evaluation Metrics
Replacement: A natural way to build image harmonization dataset is collecting a set of foreground images captured in different illumination conditions, followed by replacing one foreground with another counterpart. For example, Transient Attributes Database [84] contains 101 sets, in which each set has well-aligned images for the same scene captured in different conditions (e.g., weather, time of the day, season). This dataset has been used to construct pairs in Hday2night sub-dataset in iHarmony4 [25]. However, collecting the dataset like [84] calls for capturing the same scene with a fixed camera for a long time, which is hard to be realized in practice. To obtain the foregrounds in different capture conditions, Song et al. [165] proposed an interesting way to construct GMS Dataset. Specifically, they place the same physical model (3D foreground object) in different lighting conditions to capture different images and align the foregrounds in different images. Nevertheless, the collection cost is still very high and the diversity of foreground is restricted. In lieu of capturing images in the real world, an easy alternative is varying the illumination condition in a virtual environment. Cao et al. [10] constructed RdHarmony dataset by varying the lighting condition of the same scene using 3D rendering techniques. Within a set of images with the same scene yet various lighting conditions, the composite images could be obtained by exchanging the foregrounds between two images. Similarly, Guo et al. [54] constructed HVIDIT dataset based on the rendered dataset [36]. [66, 65] rendered the same 3D foreground model using different illumination maps. However, the rendered images have large domain gap with real images, so the harmonization model trained on rendered images cannot be directly applied to real test images.
@@ IV-F Experiments
For painterly image harmonization, we use COCO [108] and WikiArt [131]. COCO [108] contains instance segmentation annotations for 80 object categories, while WikiArt [131] contains digital artistic paintings from different styles. We create composite images based on these two datasets, with the photographic objects from COCO and the painterly backgrounds from WikiArt. In Fig. 13, we show the harmonized results of different methods (SDEdit [126], CDC [56], DIB [241], DPH [124], PHDNet [11], PHDiffusion [119]). We split COCO and WikiArt into training set and test set, based on which all methods are trained and evaluated. It can be seen that the methods (DPH, PHDNet, PHDiffusion) specifically designed for painterly image harmonization significantly outperform the other methods. In the challenging cases, where the background has dense textures or abstract styles, PHDiffusion achieves remarkable performance, probably owning to the generative ability of diffusion model and the rich prior knowledge in foundation model.
@@ V-A Rendering based Methods
The traditional methods [74, 75, 110, 106] usually use rendering techniques to generate shadow for the inserted foreground object, which needs to collect or estimate the scene geometry, foreground object geometry, and scene illumination. For example, [74] proposed to collect the rough geometry information and lighting information from users, based on which rendering techniques could be employed. However, it is very tedious and sometimes impossible to collect all the required information. In [75, 110, 106], they attempted to estimate the missing information (e.g., scene geometry, lighting information) automatically. With the recovered information, the local region to place the inserted 3D object is rendered with and without the inserted foreground object. The difference between these two rendered images reveals the impact of foreground object on the background, which is added to the input composite image to produce the target image with foreground shadow. Geometry estimation and lighting estimation based on a single image have been long studied, and many different technical approaches have been developed [110, 81]. More recently, some methods [106, 44, 238, 45, 61, 206] endeavored to estimate illumination condition and scene geometry based on a single image using deep learning models, which could achieve better performance than traditional estimation models. Some works [159, 160] proposed to forecast essential geometry information (e.g., pixel height) which cooperates with user-specified illumination to render realistic shadows.
@@ V-C Datasets and Evaluation Metrics
Similar to image harmonization in Section IV, composite images without foreground shadows can be easily obtained. Nonetheless, it is very difficult to obtain paired data, i.e., a composite image without foreground shadow and a ground-truth image with foreground shadow, which are required by supervised deep learning methods on shadow generation [248, 111, 62]. Some works [248, 111] construct rendered datasets with paired data by inserting a virtual object into 3D scene and generating shadow for this object with rendering technique. ARShadowGAN [111] released a rendered dataset named Shadow-AR by inserting a foreground object into real background image and generating its corresponding shadow with rendering technique. Shadow-AR dataset contains $3,000$ quintuples, in which each quintuple consists of a composite image without foreground shadow, its corresponding ground-truth image with foreground shadow, foreground object mask, background object mask, and background shadow mask. Shadow-AR dataset only uses 13 foreground objects from ShapeNet [14] and Stanford 3D scanning repository, so the diversity of dataset is very limited. Some examples in Shadow-AR dataset are exhibited in the first row in Fig. 16, in which we show the composite image without foreground shadow, foreground object mask, and ground-truth image with foreground shadow. Similar to ARShadowGAN [111], ShadowGAN [248] also adopted rendering technique to construct a rendered dataset, which uses $9,265$ foreground objects from ShapeNet [14] and $110$ background textures (e.g., woollen, stone, tablecloth) collected from Internet. Tao et al. [179] contributed a large-scale rendering dataset called RdSOBA, which has 788 3D foreground objects and nearly 280,000 object-shadow pairs. In particular, they place a group of 3D objects in the 3D scene, and get the images without or with object shadows using rendering techniques.
@@
Although it is feasible to generate paired data using rendering technique, the rendered images have large domain gap with real images. When applying the model trained on rendered images to real images, the performances are usually significantly degraded. To overcome this drawback, Hong et al. [62] constructed paired data by manually removing the foreground shadows from real shadow images in SOBA dataset [194] to produce synthetic composite images without foreground shadows, leading to DESOBA dataset. This strategy to create synthetic composite image is similar to the backward adjustment for constructing image harmonization dataset (see Section IV). In particular, Hong et al. [62] first remove all shadows from a shadow image to create a shadow-free image. Then, one foreground shadow region in the shadow image is overlaid by the counterpart in its corresponding shadow-free image, yielding a synthetic composite image with one missing foreground shadow. DESOBA dataset contains 839 training images with totally 2,995 object-shadow pairs and 160 test images with totally 624 object-shadow pairs. Some examples in DESOBA dataset are exhibited in the second row in Fig. 16, in which we show the composite image without foreground shadow, foreground object mask, and ground-truth image with foreground shadow. As mentioned in [62], manual shadow removal is extremely expensive.
@@ V-D Experiments
We compare existing shadow generation methods ShadowGAN [248], MaskShadowGAN [64], ARShadowGAN [111], SGRNet [62], and SGDiffusion [115]. All methods are trained on the training sets of DESOBA [62] and DESOBAv2 [115], and evaluated on the test set of DESOBA. We show the shadow images generated by different methods in Fig. 15.
@@
For practical usage, foundation diffusion model (e.g, SD [152], FLUX [1]) is imperative to generate realistic shadows. More advanced foundation model usually lead to better performance, especially in the challenging cases.
@@ VI Reflection Generation
Besides shadow, reflection is another impact that the foreground object may have on the background, especially when the foreground object is above the water or on the desktop with reflective material. To address this issue, reflection generation task aims to generate plausible reflection for the foreground object to make the composite image more realistic. Similar to Section V, we divide the existing methods into rendering based methods and non-rendering based methods. ’
@@ VI-A Rendering based Methods
Ma et al. [125] first estimated the lighting information and then employed renderer to generate coarse reflection, which is further refined using neural network.
@@ VI-B Non-rendering based Methods
Some works treat reflection generation as image-to-image translation task without estimating lighting/geometry information or employing renderer. For example, the works [208, 180] employed conditional diffusion model to synthesize reflection for the inserted foreground object. The works [31, 32] adopt similar approaches but focused on mirror reflection.
@@ VI-C Datasets and Evaluation Metrics
To evaluate the quality of generated composite images with foreground reflections, Mean Square Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity index (SSIM) [154], Learned Perceptual Image Patch Similarity (LPIPS) [245] can be adopted to measure the difference between generated image and ground-truth image. User study is also necessary to ensure that generated reflections conform to human perception. Participants are asked to select the most realistic images, based on which some metrics (e.g., B-T score [7], ratio) are calculated.
@@ VI-D Experiments
We compare different methods on DEROBA dataset. The baseline is ControlNet [243] taking composite image and foreground mask as the input of control encoder. Besides, we extend ControlNet to RGDiffusion considering the property of reflection generation task. Specifically, we inject the cropped foreground into denoising UNet via cross-attention, considering that reflections are usually the horizontal or vertical mirror images of foreground objects. We show the results of ControlNet and RGDiffusion in Fig. 17, based on which RGDiffusion has more accurate shapes and clearer details.

### v8

@@ IV-A Rendering based Methods
Conventional image relighting [145, 170, 211, 251, 189] aims to adjust the appearance of an image or the object in an image as lit by novel illumination. With some adaptation, image relighting can also be used to adjust the foreground appearance according to the illumination of a new background [145, 251, 12, 70], which bears some resemblance to image harmonization. However, they usually achieve this goal by inferring explicit illumination condition, material properties, or 3D geometry, in which the supervision for these information is difficult and expensive to acquire. Besides, they generally have strong assumption about the light source, which may not generalize well to complicated real-world scenes.
@@ IV-B Non-rendering based Methods
Early traditional image harmonization methods [229, 178, 90, 173] performed color transformation on the foreground to match the low-level color statistics between foreground and background. The difference between different methods mainly lies in the matching details. For example, Xue et al. [229] predicted the histogram zone (e.g., low, middle, high) which can be best matched between foreground and background, and then adjusted the foreground color to match the selected zone between foreground and background. [178] explored decomposing an image into a multi-resolution pyramid with multiple subbands, and performing histogram matching for each subband between foreground and background. Lalonde and Efros [90] proposed to represent foreground and background with color clusters, followed by matching foreground and background color clusters. Song et al. [173] proposed to calculate the color transformation (channel-wise scales) based on the gray pixels of foreground and background, because normalized illumination color can be directly derived from the pixel values of gray pixels. In some works on sky replacement [274, 184, 192], they attempted to match the color statistics (e.g., mean, variance) between sky region and non-sky region. Broadly speaking, traditional color transfer methods [156, 222, 44, 151, 42, 2] can also be used for color matching between foreground and background to produce a harmonized image.
@@
With the emergence of image harmonization datasets consisting of paired training data (see Section IV-E), abundant image harmonization methods [193, 28, 30, 123, 129, 181, 18, 235, 210, 149, 198, 96, 261] using paired supervision have been developed. Tsai et al. [193] proposed the first end-to-end CNN network for image harmonization and leveraged auxiliary semantic segmentation branch to enhance the basic image harmonization network. Another work Sofiiuk et al. [172] also utilized high-level semantic features, which are inserted into the encoder to provide auxiliary information. Cun and Pun [32] designed an additional Spatial-Separated Attention Module to deal with foreground and background feature maps separately. Hao et al. [63] employed self-attention [206] mechanism to propagate relevant information from background to foreground. By treating different capture conditions as different domains, Cong et al. [28] proposed a domain verification discriminator to pull close the foreground domain and background domain. Similarly, Cong et al. [29] formulated image harmonization as background-guided domain translation task, in which the domain code of background is directly used to guide the harmonization process. One byproduct of [29] is predicting the inharmony level of an image by comparing the domain codes of foreground and background, so that we can selectively harmonize those inharmonious composite images. Inspired by [29], Valanarasu et al. [194] proposed to extract style code from part of background.
@@
Some works [141, 165, 16] concurred that dynamic kernels applied to feature maps can boost the harmonization performance. Furthermore, [141, 165] pointed out the importance of global information in dynamic kernel prediction. Niu et al. [140] studied domain adaptive image harmonization by treating different datasets as different domains. Specifically, an automatic augmentation network was developed to enrich the illumination diversity of a target domain with limited data.
@@ IV-C Variants of Image Harmonization Task
Subsequently, some works [110, 111, 220, 219, 257, 169, 22] focused on inharmonious region localization task, which aims to localize the inharmonious region in an image. Liang et al. [110] explored aggregating multi-scale contextual information and suppressing redundant information. The methods [111, 220] proposed to magnify the domain discrepancy between foreground and background using color mapping for ease of identifying the inharmonious region. Chen et al. [22] explored multi-view representation including frequency view and flipped view.
@@ IV-D Related Research Fields
Image harmonization is closely related to style transfer. Note that both artistic style transfer [50, 73, 147] and photorealistic style transfer [130, 107] belong to style transfer. Image harmonization is closer to photorealistic style transfer, which transfers the style of a reference photo to another input photo. There are two main differences between image harmonization and photorealistic style transfer. 1) Firstly, image harmonization adjusts the foreground appearance according to the background, which must take the foreground location into consideration due to the locality property. In contrast, photorealistic style transfer adjusts the appearance of a whole input image according to another whole reference image. 2) Secondly, the definition of “style” in photorealistic style transfer is unclear and coarsely depends on the employed style loss (e.g., Gram matrix loss [50], AdaIn loss [73]). Differently, the goal of image harmonization is clearly adjusting the illumination statistics of foreground, so that the resultant foreground looks like the same object captured in the background illumination condition.
@@ IV-E Datasets and Evaluation Metrics
Replacement: A natural way to build image harmonization dataset is collecting a set of foreground images captured in different illumination conditions, followed by replacing one foreground with another counterpart. For example, Transient Attributes Database [89] contains 101 sets, in which each set has well-aligned images for the same scene captured in different conditions (e.g., weather, time of the day, season). This dataset has been used to construct pairs in Hday2night sub-dataset in iHarmony4 [28]. However, collecting the dataset like [89] calls for capturing the same scene with a fixed camera for a long time, which is hard to be realized in practice. To obtain the foregrounds in different capture conditions, Song et al. [173] proposed an interesting way to construct GMS Dataset. Specifically, they place the same physical model (3D foreground object) in different lighting conditions to capture different images and align the foregrounds in different images. Nevertheless, the collection cost is still very high and the diversity of foreground is restricted. In lieu of capturing images in the real world, an easy alternative is to vary the illumination condition in a virtual environment. Cao et al. [10] constructed RdHarmony dataset by varying the lighting condition of the same scene using 3D rendering techniques. Within a set of images with the same scene yet various lighting conditions, the composite images could be obtained by exchanging the foregrounds between two images. Similarly, Guo et al. [59] constructed HVIDIT dataset based on the rendered dataset [39]. [71, 70] rendered the same 3D foreground model using different illumination maps. However, the rendered images have a large domain gap with real images, so the harmonization model trained on rendered images cannot be directly applied to real test images.
@@ IV-F Experiments
For painterly image harmonization, we use COCO [115] and WikiArt [138]. COCO [115] contains instance segmentation annotations for 80 object categories, while WikiArt [138] contains digital artistic paintings from different styles. We create composite images based on these two datasets, with the photographic objects from COCO and the painterly backgrounds from WikiArt. In Fig. 13, we show the harmonized results of different methods (SDEdit [133], CDC [61], DIB [250], DPH [131], PHDNet [11], PHDiffusion [126]). We split COCO and WikiArt into training set and test set, based on which all methods are trained and evaluated. It can be seen that the methods (DPH, PHDNet, PHDiffusion) specifically designed for painterly image harmonization significantly outperform the other methods. In the challenging cases, where the background has dense textures or abstract styles, PHDiffusion achieves remarkable performance, probably due to the generative ability of diffusion model and the rich prior knowledge in foundation model.
@@ V-A Rendering based Methods
The traditional methods [79, 80, 117, 113] usually use rendering techniques to generate shadow for the inserted foreground object, which need to collect or estimate the scene geometry, foreground object geometry, and scene illumination. For example, [79] proposed to collect the rough geometry information and lighting information from users, based on which rendering techniques could be employed. However, it is very tedious and sometimes impossible to collect all the required information. In [80, 117, 113], they attempted to estimate the missing information (e.g., scene geometry, lighting information) automatically. With the recovered information, the local region to place the inserted 3D object is rendered with and without the inserted foreground object. The difference between these two rendered images reveals the impact of foreground object on the background, which is added to the input composite image to produce the target image with foreground shadow. Geometry estimation and lighting estimation based on a single image have been long studied, and many different technical approaches have been developed [117, 86]. More recently, some methods [113, 48, 247, 49, 66, 214] endeavored to estimate illumination condition and scene geometry based on a single image using deep learning models, which could achieve better performance than traditional estimation models. Some works [167, 168] proposed to forecast essential geometry information (e.g., pixel height) which cooperates with user-specified illumination to render realistic shadows.
@@ V-C Datasets and Evaluation Metrics
Similar to image harmonization in Section IV, composite images without foreground shadows can be easily obtained. Nonetheless, it is very difficult to obtain paired data, i.e., a composite image without foreground shadow and a ground-truth image with foreground shadow, which are required by supervised deep learning methods on shadow generation [258, 118, 67]. Some works [258, 118] construct rendered datasets with paired data by inserting a virtual object into 3D scene and generating shadow for this object with rendering technique. ARShadowGAN [118] released a rendered dataset named Shadow-AR by inserting a foreground object into real background image and generating its corresponding shadow with rendering technique. Shadow-AR dataset contains $3,000$ quintuples, in which each quintuple consists of a composite image without foreground shadow, its corresponding ground-truth image with foreground shadow, foreground object mask, background object mask, and background shadow mask. Shadow-AR dataset only uses 13 foreground objects from ShapeNet [14] and Stanford 3D scanning repository, so the diversity of dataset is very limited. Some examples in Shadow-AR dataset are exhibited in the first row in Fig. 16, in which we show the composite image without foreground shadow, foreground object mask, and ground-truth image with foreground shadow. Similar to ARShadowGAN [118], ShadowGAN [258] also adopted rendering technique to construct a rendered dataset, which uses $9,265$ foreground objects from ShapeNet [14] and $110$ background textures (e.g., woolen, stone, tablecloth) collected from Internet. Tao et al. [187] contributed a large-scale rendering dataset called RdSOBA, which has 788 3D foreground objects and nearly 280,000 object-shadow pairs. In particular, they place a group of 3D objects in the 3D scene, and get the images without or with object shadows using rendering techniques.
@@
Although it is feasible to generate paired data using rendering technique, the rendered images have large domain gap with real images. When applying the model trained on rendered images to real images, the performances are usually significantly degraded. To overcome this drawback, Hong et al. [67] constructed paired data by manually removing the foreground shadows from real shadow images in SOBA dataset [202] to produce synthetic composite images without foreground shadows, leading to DESOBA dataset. This strategy to create synthetic composite images is similar to the backward adjustment for constructing image harmonization dataset (see Section IV). In particular, Hong et al. [67] first remove all shadows from a shadow image to create a shadow-free image. Then, one foreground shadow region in the shadow image is overlaid by the counterpart in its corresponding shadow-free image, yielding a synthetic composite image with one missing foreground shadow. DESOBA dataset contains 839 training images with totally 2,995 object-shadow pairs and 160 test images with totally 624 object-shadow pairs. Some examples in DESOBA dataset are exhibited in the second row in Fig. 16, in which we show the composite image without foreground shadow, foreground object mask, and ground-truth image with foreground shadow. As mentioned in [67], manual shadow removal is extremely expensive.
@@ V-D Experiments
We compare existing shadow generation methods ShadowGAN [258], MaskShadowGAN [69], ARShadowGAN [118], SGRNet [67], and SGDiffusion [122]. All methods are trained on the training set of DESOBA [67] and DESOBAv2 [122], and evaluated on the test set of DESOBA. We show the shadow images generated by different methods in Fig. 15.
@@
For practical usage, foundation diffusion model (e.g., SD [160], FLUX [1]) is imperative to generate realistic shadows. More advanced foundation models usually lead to better performance, especially in the challenging cases.
@@ VI Reflection Generation
Besides shadow, reflection is another impact that the foreground object may cast on the background, especially when the foreground object is above the water or on the desktop with reflective material. To address this issue, reflection generation task aims to generate plausible reflection for the foreground object to make the composite image more realistic. Similar to Section V, we divide the existing methods into rendering based methods and non-rendering based methods. ’
@@ VI-A Rendering based Methods
Ma et al. [132] first estimated the lighting information and then employed a renderer to generate coarse reflection, which is further refined using neural network.
@@ VI-B Non-rendering based Methods
Some works treat reflection generation as image-to-image translation task without estimating lighting/geometry information or employing renderer. For example, the works [216, 188] employed conditional diffusion model to synthesize reflection for the inserted foreground object. The works [34, 35] adopted similar approaches but focused on mirror reflection.
@@ VI-C Datasets and Evaluation Metrics
To evaluate the quality of generated composite images with foreground reflections, Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity index (SSIM) [162], Learned Perceptual Image Patch Similarity (LPIPS) [255] can be adopted to measure the difference between generated image and ground-truth image. User study is also necessary to ensure that generated reflections conform to human perception. Participants are asked to select the most realistic images, based on which some metrics (e.g., B-T score [7], ratio) are calculated.
@@ VI-D Experiments
We compare different methods on DEROBA dataset. The baseline is ControlNet [252] taking composite image and foreground mask as the input of control encoder. Besides, we extend ControlNet to RGDiffusion considering the property of reflection generation task. Specifically, we inject the cropped foreground into denoising UNet via cross-attention, considering that reflections are usually the horizontal or vertical mirror images of foreground objects. As shown in Fig. 17, RGDiffusion produces more accurate shapes and clearer details compared with ControlNet.
For practical usage, diffusion-based methods like RGDiffusion can usually produce satisfactory results.